the sum of two numbers is 85. Two times the smaller number exceeds the larger number by26. Find the numbers
Answer by rothauserc(4718)

let x be the smaller number and y the other number, then we have
x + y = 85
2x = y +26
solve second equation for x
x = (y +26)/2
substitute for x in first equation
(y +26) / 2 + y = 85
multiply both sides of = by 2
y + 26 +2y = 170
combine like terms
3y = 144
y = 48
now x = 85 -48 = 37
x = 37 and y = 48
check answer by substituting for x and y in second equation
2*37 = 48 +26
74 = 74
our answer checks :-)